OpenStudy (anonymous):

which is larger 0.23 or 1/4 why? I think they are the same

OpenStudy (baldymcgee6):

well 1/4 = 0.25

OpenStudy (unklerhaukus):

nope, 1/4 = 25/100=0.25

OpenStudy (baldymcgee6):

\[0.25 \neq 0.23\]

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

0.25>0.23

OpenStudy (unklerhaukus):

0.23=23/100

OpenStudy (anonymous):

right..okay I was unsure..thanks guys